OpenStudy (yamyam70):

Every real number is raised to the 1st power, Agree / Disagree ?

OpenStudy (anonymous):

I don't know what you mean by that.

OpenStudy (yamyam70):

Example 1,2,3,4,5 .... all of them has the exponent 1

OpenStudy (anonymous):

Raising a number to a power is an operation you perform on a number, not an inherent property of the number itself. It's certainly true that 1^1 = 1, 2^2 = 2, 3^1 = 3, etc. but can't say that the number 3 all by itself is not being raised to a power.

OpenStudy (yamyam70):

Thanks for clearing that up @Jemurray3 :))

OpenStudy (anonymous):

I'm sorry, that should be *you can't say that the number 3 all by itself is being raised to a power"

OpenStudy (anonymous):

It doesn't make much difference in the case of raising something to the power one, because that doesn't change anything. But no problem.

OpenStudy (yamyam70):

I see :)